OpenStudy (anonymous):

2r2 + 3s3 - r2 + 4t2 - r2 could someone please explain how to find the answer using r = -2, s = -3, and t = 5. every time I do it, and submit my answer the question is always wrong.

OpenStudy (anonymous):

Use the proper notation: 2r^2 + 3s^3 - r^2 + 4t^2 - r^2 = 2(-2)^2 + 3(-3)^3 - (-2)^2 + 4(5^2) - (-2)^2 = 2(4) + 3(-27) - 4 + 100 - 4 = 8 - 81 - 4 + 100 - 4 = 4 + 19 - 4 = 19

OpenStudy (anonymous):

Thank you so much! Now I think I'm starting to understand why I wasn't getting it right.

OpenStudy (anonymous):

You're welcome!! Xo!

OpenStudy (anonymous):

If you need more help, then come by me and tell me!! Xo!